OpenStudy (123ab456c):

A square pyramid has a base area of 25 ft2. The pyramid’s total surface area is 105 ft2. What is the area of each triangular face? A. 20 ft2 B. 26 ft2 C. 40 ft2 D. 80 ft2

jimthompson5910 (jim_thompson5910):

how many faces does a square pyramid have?

OpenStudy (123ab456c):

@jim_thompson5910 do you know the answer

jimthompson5910 (jim_thompson5910):

here's what a square pyramid looks like http://towerbridgeproject10.wikispaces.com/file/view/square_pyramid.JPG/136947183/269x279/square_pyramid.JPG

OpenStudy (123ab456c):

it has 4 faces

jimthompson5910 (jim_thompson5910):

5 faces actually 4 lateral faces (that are triangles) and one square face on the bottom

OpenStudy (anonymous):

So first step is to subtract the area of the base from the total surface area because the rest of the faces are the same

OpenStudy (anonymous):

Which is...?

jimthompson5910 (jim_thompson5910):

B = area of base (square base) L = area of each triangle there is 1 base and 4 triangles total surface area = B + 4*L

OpenStudy (123ab456c):

is the answer C

OpenStudy (anonymous):

No it isn't try using the equation jim gave

jimthompson5910 (jim_thompson5910):

total surface area = 105 area of base = B = 25 so, total surface area = B + 4*L 105 = 25 + 4*L solve for L

OpenStudy (anonymous):

\[25x+zy=105\] x= number of square faces y=number of triangular faces

OpenStudy (123ab456c):

can you tell me the answer and then explain it that way i wont get confused

OpenStudy (anonymous):

Alright it's A

OpenStudy (anonymous):

So there is a total of 105 square units 25 square units is the base And there are four identical triangular faces with unknown area

OpenStudy (anonymous):

So try solving the equation I gave

OpenStudy (123ab456c):

i did it makes SENSE KNOW thank you so much now i understand it !!!:):)

OpenStudy (anonymous):

Youre welcome :D